# Generalized Linear Bandits with Limited Adaptivity

### v4

@@ 4 RS-GLinCB
1:  Initialize: $\mathbf{V}={\mathbf{H}}_{1}=\lambda{\mathbf{I}}$, ${\mathcal{T}}_{o}=\emptyset$, $\tau=1$, $\lambda\coloneqq d\log({T/\delta})/R^{2}$ and $\gamma\coloneqq 25RS\sqrt{d\log\left(\frac{T}{\delta}\right)}$.
@@
7:        Compute $\widehat{\theta}_{o}=\operatorname*{arg\,min}_{\theta}\sum_{s\in{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$.
8:     else
@@
10:           Set $\tau=t$ and $\widetilde{\theta}\leftarrow\operatorname*{arg\,min}_{\theta}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$ and
11:           $\widehat{\theta}_{\tau}\leftarrow\texttt{Project}(\widetilde{\theta})$.
12:        end if
13:        Update ${\mathcal{X}}_{t}\leftarrow{\mathcal{X}}_{t}\setminus\{x\in{\mathcal{X}}_{t}:UCB_{o}(x)<\max_{z\in{\mathcal{X}}_{t}}LCB_{o}(z)\}$.
@@
14:        Select $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})$ and observe reward $r_{t}$.
@@
15:        Update ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+\frac{\dot{\mu}\left(\langle x_{t},\widehat{\theta}_{o}\rangle\right)}{e}x_{t}x_{t}^{\intercal}$.
16:     end if
17:  end for
@@
In this section we present RS-GLinCB (Algorithm 2) that solves Problem 2 described in Section 2, which enforces constraints of limited adaptivity setting M2. This algorithm incorporates a novel switching criterion (Line 4), extending the determinant-doubling approach of Abbasi-Yadkori et al. (2011). Additionally, we introduce an arm-elimination step (Line 13) to obtain tighter regret guarantees. Throughout this section, we set $\lambda=d\log({T/\delta})/R^{2}$ and $\gamma=25RS\sqrt{d\log\left({T}/{\delta}\right)}$.
@@
At round $t$, on receiving an arm set ${\mathcal{X}}_{t}$, RS-GLinCB first checks the Switching Criterion I (Line 4). This criterion checks whether for any arm $x\in{\mathcal{X}}_{t}$ the quantity $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ is greater than a carefully chosen $\kappa$-dependent threshold. Here $\mathbf{V}$ is the design matrix corresponding to all arms that have been played in the rounds in ${\mathcal{T}}_{o}$ ($\coloneqq$ the set of rounds preceding round $t$, where Switching Criterion I was triggered). Under this criterion the arm that maximizes $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ is played (call this arm $x_{t}$) and the corresponding reward is obtained. Subsequently in Line 6, the set ${\mathcal{T}}_{o}$ is updated to include $t$; the design matrix $\mathbf{V}$ is updated as $\mathbf{V}\leftarrow\mathbf{V}+x_{t}x_{t}^{\intercal}$; and the scaled design matrix ${\mathbf{H}}_{t+1}$ is set to ${\mathbf{H}}_{t}$. The MLE is computed (Line 7) based on the data in the rounds in ${\mathcal{T}}_{o}$ to obtain $\widehat{\theta}_{o}$.
When Switching Criterion I is not triggered, the algorithm first checks (Line 9) the Switching Criterion II, that is whether the determinant of the scaled design matrix ${\mathbf{H}}_{t}$ has become more than double of that of ${\mathbf{H}}_{\tau}$ (where $\tau$ is the last round before $t$ when Switching Criterion II was triggered). If Switching Criterion II is triggered at round $t$, then in Line 10, the algorithm sets $\tau\leftarrow t$ and recomputes the MLE over all the past rounds except those in ${\mathcal{T}}_{o}$ to obtain $\widetilde{\theta}$. Then $\widetilde{\theta}$ is projected into an ellipsoid around $\widehat{\theta}_{o}$ to obtain the estimate $\widehat{\theta}_{\tau}$ via the following optimization problem [^9] This optimization problem is non-convex. However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$.,
@@
$\min_{\theta}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu(\langle x_{s},\widetilde{\theta}\rangle)\right)x_{s}\right\rVert_{{\mathbf{H}}(\theta)}\text{s.t. }\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}.$ (8)
@@
Here ${\mathbf{H}}(\theta)\coloneqq\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\mathsf{T}}$. After checking Switching Criterion II, the algorithm performs an arm elimination step (Line 13) based on the parameter estimate $\widehat{\theta}_{o}$ as follows: for every arm $x\in{\mathcal{X}}_{t}$, we compute $UCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and $LCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ [^10] We note that in case $\kappa$ is unknown, any known upper bound on $\kappa$ suffices for the algorithm.. Then, ${\mathcal{X}}_{t}$ is updated by eliminating from it the arms with $UCB_{o}(\cdot)$ less than the highest $LCB_{o}(\cdot)$. For arms in the reduced arm set ${\mathcal{X}}_{t}$, RS-GLinCB computes the index $UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})\coloneqq\langle x,\widehat{\theta}_{\tau}\rangle+150\left\lVert x\right\rVert_{{\mathbf{H}}_{\tau}^{-1}}\sqrt{d\log{({T}/{\delta})}}$, and plays the arm $x_{t}$ with the highest index (Line 14). After observing the subsequent reward $r_{t}$, the algorithm updates the scaled design matrix ${\mathbf{H}}_{t}$ (Line 15) as follows: ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+(\dot{\mu}(\langle x_{t},\widehat{\theta}_{o}\rangle)/e)x_{t}x_{t}^{\intercal}$. With this, the round $t$ ends and the algorithm moves to the next round. Next, in Lemma 4.1 and Theorem 4.2 we present the guarantees on number of policy updates and regret, respectively, for RS-GLinCB. Detailed proofs for both are provided in Appendix B.
@@
RS-GLinCB (Algorithm 2), during its entire execution, updates its policy at most $O(R^{4}S^{2}\ \kappa d^{2}\ \log^{2}(T/\delta))$ times.
@@
Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}d\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{RT}/{\delta}}\right)+\ \kappa d^{2}R^{5}S^{2}\log^{2}\left({{T}/{\delta}}\right)\big{)}.$
@@
Switching Criterion I is essential in delivering tight regret guarantees in the non-linear setting. Unlike existing literature Faury et al. (2022), which relies on warm-up rounds based on observed rewards (hence heavily dependent on reward models), RS-GLinCB presents a context-dependent criterion that implicitly checks whether the estimate $\dot{\mu}(\langle x,\widehat{\theta}_{o}\rangle)$ is within a constant factor of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemmas B.3 and B.4). We show that the number of times Switching Criterion I is triggered is only $O(\kappa d^{2}\log^{2}(T))$ (see Lemma B.11), hence incurring a small regret in these rounds.
@@
Unlike Abbasi-Yadkori et al. (2011), our determinant-doubling Switching Criterion II uses the scaled design matrix ${\mathbf{H}}_{t}$ instead of the unscaled version (similar to $\mathbf{V}$). The matrix ${\mathbf{H}}_{t}$, estimating the Hessian of the log-loss, is crucial for achieving optimal regret. This modification is crucial in extending algorithms satisfying limited adaptivity setting M2 for the CB problem with a linear reward model to more general GLM reward models.
@@
The feasible set for the optimization stated in 8 is an ellipsoid around $\widehat{\theta}_{o}$, which contains $\theta^{*}$ with high probability. Deviating from existing literature on GLM Bandits which projects the estimate into the ball set of radius $S$ ($\{\theta:\left\lVert\theta\right\rVert\leq S\}$), our projection step leads to tighter regret guarantees; notably, the leading $\sqrt{T}$ term is free of parameters $S$ (and $R$). This resolves the conjecture made in Lee et al. (2024) regarding the possibility of obtaining $S$-free regret in the $\sqrt{T}$ term in logistic bandits.
Remark 4.6.
The regret guarantees of the logistic bandit algorithms in Abeille et al. (2021); Lee et al. (2024) have a second-order term that is minimum of an arm-geometry dependent quantity (see Theorem 3 of Lee et al. (2024)) and a $\kappa$-dependent term similar to our regret guarantee. Although our analysis is not able to accommodate this arm-geometry dependent quantity, we underscore that our algorithm is computationally efficient while the above works are not. In fact, to the best of our knowledge, the other known efficient algorithms for logistic bandits Faury et al. (2022); Zhang and Sugiyama (2023) also do not achieve the arm-geometry dependent regret term. It can be interesting to design an efficient algorithm that is able to achieve the same guarantees in the second-order regret term as in Abeille et al. (2021); Lee et al. (2024).
@@ 5 Experiments
Logistic. We compared RS-GLinCB against ECOLog Faury et al. (2022) and GLOC Jun et al. (2017), the only algorithms with overall time complexity $\tilde{O}(T)$ for this setting. The dimension was set to $d=5$, number of arms per round to $K=20$, and $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=5$. Arms were sampled uniformly from the $d$-dimensional unit ball. We ran simulations for $T=20,000$ rounds, repeating them 10 times. RS-GLinCB showed the smallest regret with a flattened regret curve, as seen in Fig. 1 (top-left).
Probit. For the probit reward model, we compared RS-GLinCB against GLOC and GLM-UCB Filippi et al. (2010). The dimension was set to $d=5$ and number of arms per round to $K=20$. $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=3$. Arm features were generated similarly as in the logistic bandit simulation. We ran simulations for $T=5,000$ rounds, repeating them 10 times. RS-GLinCB outperformed both baselines, as shown in Fig. 1 (top-right).
@@ 6 Conclusion and Future Work
The Contextual Bandit problem with GLM rewards is a ubiquitous framework for studying online decision-making with non-linear rewards. We study this problem with a focus on limited adaptivity. In particular, we design algorithms B-GLinCB and RS-GLinCB that obtain optimal regret guarantees for two prevalant limited adaptivity settings M1 and M2 respectively. A key feature of our guarantees are that their leading terms are independent of an instance dependent parameter $\kappa$ that captures non-linearity. To the best of our knowledge, our paper provides the first algorithms for the CB problem with GLM rewards under limited adaptivity (and otherwise) that achieve $\kappa$-independent regret. The regret guarantee of RS-GLinCB, not only aligns with the best-known guarantees for Logistic Bandits but enhances them by removing the dependence on $S$ (upper bound on $\left\lVert\theta^{*}\right\rVert$) in the leading term of the regret and therefore resolves a conjecture in Lee et al. (2024). The batch learning algorithm B-GLinCB, for $M=\Omega(\log{(\log{T})})$, achieves a regret of $\widetilde{O}\left(dRS\left(\sqrt{{d/{\widehat{\kappa}}}}\land\sqrt{1/\kappa^{*}}\right)\sqrt{T}\right)$. We believe that the dependence on $d$ along with the ${\widehat{\kappa}}$ term is not tight and improving the dependence is a relevant direction for future work.

### v5

@@ 4 RS-GLinCB
1:  Initialize: $\mathbf{V}={\mathbf{H}}_{1}=\lambda{\mathbf{I}}$, ${\mathcal{T}}_{o}=\emptyset$, $\tau=1$, $\lambda\coloneqq dR^{2}\log({T/\delta})/S$ and $\gamma\coloneqq 25R^{2}S^{\nicefrac{{3}}{{2}}}\sqrt{d\log\left(\frac{ST}{\delta}\right)}$.
@@
7:        Compute $\widetilde{\theta}_{o}=\operatorname*{arg\,min}_{\theta}\sum_{s\in{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$.
8:        $\widehat{\theta}_{o}\leftarrow\texttt{Project}(\widetilde{\theta}_{o})$ via (8).
9:     else
@@
11:           Set $\tau=t$ and $\widetilde{\theta}\leftarrow\operatorname*{arg\,min}_{\theta}\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\ell(\theta,x_{s},r_{s})+\frac{\lambda}{2}\left\lVert\theta\right\rVert_{2}^{2}$ and
12:           $\widehat{\theta}_{\tau}\leftarrow\texttt{Project}(\widetilde{\theta})$ via (9).
13:        end if
14:        Update ${\mathcal{X}}_{t}\leftarrow{\mathcal{X}}_{t}\setminus\{x\in{\mathcal{X}}_{t}:UCB_{o}(x)<\max_{z\in{\mathcal{X}}_{t}}LCB_{o}(z)\}$.
@@
15:        Select $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})$ and observe reward $r_{t}$.
@@
16:        Update ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+\frac{\dot{\mu}\left(\langle x_{t},\widehat{\theta}_{o}\rangle\right)}{e}x_{t}x_{t}^{\intercal}$.
17:     end if
18:  end for
@@
In this section we present RS-GLinCB (Algorithm 2) that solves Problem 2 described in Section 2, which enforces constraints of limited adaptivity setting M2. This algorithm incorporates a novel switching criterion (Line 4), extending the determinant-doubling approach of AYPS (11). Additionally, we introduce an arm-elimination step (Line 14) to obtain tighter regret guarantees. Throughout this section, we set $\lambda=dR^{2}\log({T/\delta})/S$ and $\gamma=25R^{2}S^{\nicefrac{{3}}{{2}}}\sqrt{d\log\left({ST}/{\delta}\right)}$.
@@
At round $t$, on receiving an arm set ${\mathcal{X}}_{t}$, RS-GLinCB first checks the Switching Criterion I (Line 4). This criterion checks whether for any arm $x\in{\mathcal{X}}_{t}$ the quantity $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ is greater than a carefully chosen $\kappa$-dependent threshold. Here $\mathbf{V}$ is the design matrix corresponding to all arms that have been played in the rounds in ${\mathcal{T}}_{o}$ ($\coloneqq$ the set of rounds preceding round $t$, where Switching Criterion I was triggered). Under this criterion the arm that maximizes $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ is played (call this arm $x_{t}$) and the corresponding reward is obtained. Subsequently in Line 6, the set ${\mathcal{T}}_{o}$ is updated to include $t$; the design matrix $\mathbf{V}$ is updated as $\mathbf{V}\leftarrow\mathbf{V}+x_{t}x_{t}^{\intercal}$; and the scaled design matrix ${\mathbf{H}}_{t+1}$ is set to ${\mathbf{H}}_{t}$. The MLE is computed (Line 7) based on the data in the rounds in ${\mathcal{T}}_{o}$ to obtain $\widetilde{\theta}_{o}$. This is subsequently projected to the ball $\{\theta\in\mathbb{R}^{d}:\left\lVert\theta\right\rVert\leq S\}$ via the following optimization problem [^8] This optimization problem is non-convex. However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$.,
@@
$\widehat{\theta}_{o}=\operatorname*{arg\,min}_{\left\lVert\theta\right\rVert\leq S}\left\lVert\sum_{s\in{\mathcal{T}}_{o}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu\left(\langle x_{s},\widetilde{\theta}_{o}\rangle\right)\right)-\lambda\widetilde{\theta}_{o}\right\rVert_{{\mathbf{H}}_{w}(\theta)^{-1}}$ (8)
where ${\mathbf{H}}_{w}(\theta)\coloneqq\sum_{s\in{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$.
When Switching Criterion I is not triggered, the algorithm first checks (Line 10) the Switching Criterion II, that is whether the determinant of the scaled design matrix ${\mathbf{H}}_{t}$ has become more than double of that of ${\mathbf{H}}_{\tau}$ (where $\tau$ is the last round before $t$ when Switching Criterion II was triggered). If Switching Criterion II is triggered at round $t$, then in Line 11, the algorithm sets $\tau\leftarrow t$ and recomputes the MLE over all the past rounds except those in ${\mathcal{T}}_{o}$ to obtain $\widetilde{\theta}$. Then $\widetilde{\theta}$ is projected into an ellipsoid around $\widehat{\theta}_{o}$ to obtain the estimate $\widehat{\theta}_{\tau}$ via the following optimization problem [^9] This optimization problem is non-convex. However, a convex relation of this optimization problem is detailed in Appendix E, which leads to slightly worse regret guarantees in $\texttt{poly}(R,S)$.,
$\widehat{\theta}_{\tau}=\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{d}}\left\lVert\sum_{s\in{[t-1]\setminus{\mathcal{T}}_{o}}}\left(\mu\left(\langle x_{s},\theta\rangle\right)-\mu(\langle x_{s},\widetilde{\theta}\rangle)\right)x_{s}-\lambda\widetilde{\theta}\right\rVert_{{\mathbf{H}}(\theta)^{-1}}\text{s.t. }\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}.$ (9)
@@
Here ${\mathbf{H}}(\theta)\coloneqq\sum_{s\in[t-1]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x_{s},\theta\rangle\right)x_{s}x_{s}^{\mathsf{T}}$. After checking Switching Criterion II, the algorithm performs an arm elimination step (Line 14) based on the parameter estimate $\widehat{\theta}_{o}$ as follows: for every arm $x\in{\mathcal{X}}_{t}$, we compute $UCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ and $LCB_{o}(x)=\langle x,\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$ [^10] We note that in case $\kappa$ is unknown, any known upper bound on $\kappa$ suffices for the algorithm.. Then, ${\mathcal{X}}_{t}$ is updated by eliminating from it the arms with $UCB_{o}(\cdot)$ less than the highest $LCB_{o}(\cdot)$. For arms in the reduced arm set ${\mathcal{X}}_{t}$, RS-GLinCB computes the index $UCB(x,{\mathbf{H}}_{\tau},\widehat{\theta}_{\tau})\coloneqq\langle x,\widehat{\theta}_{\tau}\rangle+150\left\lVert x\right\rVert_{{\mathbf{H}}_{\tau}^{-1}}R\sqrt{dS\log(ST/\delta)}$, and plays the arm $x_{t}$ with the highest index (Line 15). After observing the subsequent reward $r_{t}$, the algorithm updates the scaled design matrix ${\mathbf{H}}_{t}$ (Line 16) as follows: ${\mathbf{H}}_{t+1}\leftarrow{\mathbf{H}}_{t}+(\dot{\mu}(\langle x_{t},\widehat{\theta}_{o}\rangle)/e)x_{t}x_{t}^{\intercal}$. With this, the round $t$ ends and the algorithm moves to the next round. Next, in Lemma 4.1 and Theorem 4.2 we present the guarantees on number of policy updates and regret, respectively, for RS-GLinCB. Detailed proofs for both are provided in Appendix B.
@@
RS-GLinCB (Algorithm 2), during its entire execution, updates its policy at most $O(\kappa d^{2}R^{6}S^{3}\log^{2}(T/\delta))$ times.
@@
Given $\delta\in(0,1)$, with probability $\geq 1-\delta$, the regret of RS-GLinCB (Algorithm 2) satisfies $\textsc{R}_{T}=O\big{(}dRS^{\nicefrac{{1}}{{2}}}\sqrt{\sum_{t\in[T]}\dot{\mu}\left(\langle x_{t}^{*},\theta^{*}\rangle\right)}\log\left({{ST}/{\delta}}\right)+\ \kappa d^{2}R^{7}S^{3}\log^{2}\left({{ST}/{\delta}}\right)\big{)}.$
@@
Switching Criterion I is essential in delivering tight regret guarantees in the non-linear setting. Unlike existing literature FAJC (22), which relies on warm-up rounds based on observed rewards (hence heavily dependent on reward models), RS-GLinCB presents a context-dependent criterion that implicitly checks whether the estimate $\dot{\mu}(\langle x,\widehat{\theta}_{o}\rangle)$ is within a constant factor of $\dot{\mu}\left(\langle x,\theta^{*}\rangle\right)$ (see Lemmas B.4 and B.5). We show that the number of times Switching Criterion I is triggered is only $O(\kappa d^{2}\log^{2}(T))$ (see Lemma B.12), hence incurring a small regret in these rounds.
@@
Unlike AYPS (11), our determinant-doubling Switching Criterion II uses the scaled design matrix ${\mathbf{H}}_{t}$ instead of the unscaled version (similar to $\mathbf{V}$). The matrix ${\mathbf{H}}_{t}$, estimating the Hessian of the log-loss, is crucial for achieving optimal regret. This modification is crucial in extending algorithms satisfying limited adaptivity setting M2 for the CB problem with a linear reward model to more general GLM reward models.
@@
The regret guarantees of the logistic bandit algorithms in AFC (21); LYJ (24) have a second-order term that is minimum of an arm-geometry dependent quantity (see Theorem 3 of LYJ (24)) and a $\kappa$-dependent term similar to our regret guarantee. Although our analysis is not able to accommodate this arm-geometry dependent quantity, we underscore that our algorithm is computationally efficient while the above works are not. In fact, to the best of our knowledge, the other known efficient algorithms for logistic bandits FAJC (22); ZS (23) also do not achieve the arm-geometry dependent regret term. It can be interesting to design an efficient algorithm that is able to achieve the same guarantees in the second-order regret term as in AFC (21); LYJ (24).
@@ 5 Experiments
Logistic. We compared RS-GLinCB against ECOLog FAJC (22) and GLOC JBNW (17), the only algorithms with overall time complexity $\tilde{O}(T)$ for this setting. The dimension was set to $d=5$, number of arms per round to $K=20$, and $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=5$. Arms were sampled uniformly from the $d$-dimensional unit ball. We ran simulations for $T=20,000$ rounds, repeating them 10 times. RS-GLinCB showed the smallest regret with a flattened regret curve, as seen in Fig. 1 (top-left).
Probit. For the probit reward model, we compared RS-GLinCB against GLOC and GLM-UCB FCGS (10). The dimension was set to $d=5$ and number of arms per round to $K=20$. $\theta^{*}$ was sampled from a $d$-dimensional sphere of radius $S=3$. Arm features were generated similarly as in the logistic bandit simulation. We ran simulations for $T=5,000$ rounds, repeating them 10 times. RS-GLinCB outperformed both baselines, as shown in Fig. 1 (top-right).
@@ 6 Conclusion and Future Work
The Contextual Bandit problem with GLM rewards is a ubiquitous framework for studying online decision-making with non-linear rewards. We study this problem with a focus on limited adaptivity. In particular, we design algorithms B-GLinCB and RS-GLinCB that obtain optimal regret guarantees for two prevalant limited adaptivity settings M1 and M2 respectively. A key feature of our guarantees are that their leading terms are independent of an instance dependent parameter $\kappa$ that captures non-linearity. To the best of our knowledge, our paper provides the first algorithms for the CB problem with GLM rewards under limited adaptivity (and otherwise) that achieve $\kappa$-independent regret. The regret guarantee of RS-GLinCB, aligns with the best-known guarantees for Logistic Bandits. The batch learning algorithm B-GLinCB, for $M=\Omega(\log{(\log{T})})$, achieves a regret of $\widetilde{O}\left(dRS\left(\sqrt{{d/{\widehat{\kappa}}}}\land\sqrt{1/\kappa^{*}}\right)\sqrt{T}\right)$. We believe that the dependence on $d$ along with the ${\widehat{\kappa}}$ term is not tight and improving the dependence is a relevant direction for future work.